OpenStudy (anonymous):

derivate (x^(1/3)+x)/(x^(1/2))

OpenStudy (anonymous):

find the derivative of this function, plz.

OpenStudy (anonymous):

split it up over the denominator

OpenStudy (anonymous):

its quicker than quotient rule

OpenStudy (anonymous):

ie f(x) = [ x^(1/3) / x^(1/2) ] + [ x / x^(1/2) ]

OpenStudy (anonymous):

(1/3) -(1/2) = -1/6

OpenStudy (amistre64):

(x^(1/3)+x) (x^(-1/2)) x^(1/3-1/2) + x^(1-1/2)

OpenStudy (anonymous):

therefore f(x) = x^(-1/6) + x^(1/2)

OpenStudy (anonymous):

then power rule

OpenStudy (anonymous):

f ' (x) = (-1/6) x^(-7/6) +(1/2) x^(-1/2)

OpenStudy (anonymous):

please i need the answer by hand writing. in pdf if possible.

OpenStudy (amistre64):

then the best results would be for you to pick up a pencil and paper and write down what elec has done instead wanting someone to go out of their way even more for free...... just a thought.

OpenStudy (shadowfiend):

Right on, amistre.